OpenStudy (juscallmesteve):

Would I use L'Hospitals rule to solve Lim x-->INF x^2e^-x

OpenStudy (juscallmesteve):

\[\lim_{x \rightarrow \infty} x ^{2}e ^{-x}\]

OpenStudy (abb0t):

Yeah, but you'd need to apply it more than once.

OpenStudy (juscallmesteve):

thank you

OpenStudy (welshfella):

yep

OpenStudy (juscallmesteve):

so would \[\lim_{x \rightarrow \infty} x^2e ^{-x} = 0\]

OpenStudy (abb0t):

you're really looking at \(\large \sf \frac{x^2}{e^x}\) if that helps.

OpenStudy (juscallmesteve):

how I got that is first I rewrote the problem as \[\lim_{x \rightarrow \infty} \frac{ x^2 }{ e^x }\] and that would be \[\infty \over \infty \]

OpenStudy (juscallmesteve):

so do it again \[\lim_{x \rightarrow \infty} \frac{ 2x }{ e^x } \] is \[\infty \over \infty \]

OpenStudy (abb0t):

do it again.

OpenStudy (abb0t):

L'hopitals can be applied more than once, and often is the case.

OpenStudy (juscallmesteve):

then do hospitals again \[\lim_{x \rightarrow \infty} \frac{ 2 }{ e^x }\] limit is now \[2 \over \infty \] which is like saying 0

OpenStudy (abb0t):

\[\sf \lim_{x \rightarrow ∞} \frac{1}{e^x} \]

OpenStudy (abb0t):

which is something you should know by now.

OpenStudy (juscallmesteve):

why would it be 1 over? the d/dx of (2x) is 2?

OpenStudy (abb0t):

d\dx of 2x = 2 factor out \[\sf 2 \times \lim_{x \rightarrow ∞} \frac{1}{e^x}\]

OpenStudy (abb0t):

"factor"

OpenStudy (juscallmesteve):

ok thank you